# STRelay: A Universal Spatio-Temporal Relaying Framework for Location Prediction with Future Spatiotemporal Contexts

### v1

@@ Abstract
Next location prediction is a critical task in human mobility modeling, enabling applications like travel planning and urban mobility management. Existing methods mainly rely on historical spatiotemporal trajectory data to train sequence models that directly forecast future locations. However, they often overlook the importance of the future spatiotemporal contexts, which are highly informative for the future locations. For example, knowing how much time and distance a user will travel could serve as a critical clue for predicting the user’s next location. Against this background, we propose STRelay, a universal SpatioTemporal Relaying framework explicitly modeling the future spatiotemporal context given a human trajectory, to boost the performance of different location prediction models. Specifically, STRelay models future spatiotemporal contexts in a relaying manner, which is subsequently integrated with the encoded historical representation from a base location prediction model, enabling multi-task learning by simultaneously predicting the next time interval, next moving distance interval, and finally the next location. We evaluate STRelay integrated with four state-of-the-art location prediction base models on four real-world trajectory datasets. Results demonstrate that STRelay consistently improves prediction performance across all cases by 3.19%-11.56%. Additionally, we find that the future spatiotemporal contexts are particularly helpful for entertainment-related locations and also for user groups who prefer traveling longer distances. The performance gain on such non-daily-routine activities, which often suffer from higher uncertainty, is indeed complementary to the base location prediction models that often excel at modeling regular daily routine patterns.
Introduction
With the rapid growth of location-based social media, users are increasingly inclined to share their activities on Location-Based Social Networks (LBSNs), generating rich digital footprints for human mobility studies (Yang et al. 2014). A key task in human mobility modeling is predicting a user’s next location based on the user’s historical mobility trajectory (Noulas et al. 2012), which is a fundamental building block supporting various location-based applications, such as travel planning (Zhou et al. 2021), taxi ride-sharing scheduling (Ma, Zheng, and Wolfson 2013), and user behavior patterns analyzing (Sun et al. 2020).
@@ Introduction
To leverage the spatiotemporal characteristics of trajectory data, various methods have been proposed for next Point of Interest (POI) prediction [^1] Given the context of human trajectories, we do not distinguish the two terms “location” and “POI” throughout this paper.. Traditional next POI prediction methods forecast future locations relying solely on historical spatiotemporal information. For example, sequential transitions are typically captured using Markov chain models (Rendle, Freudenthaler, and Schmidt-Thieme 2010; Cheng et al. 2013; Feng et al. 2015), while long-term dependencies are modeled through RNNs (Liu et al. 2016; Zhao et al. 2019) and attention-based architectures (Lian et al. 2020). More recently, some methods using additional context, such as historical spatiotemporal contexts (Yang et al. 2020) and social networks (Rao et al. 2022; Qin et al. 2023; Xu et al. 2023), have been proposed for further performance improvement. However, these methods overlook the time information of the predicted next location, and thus always make the same prediction given a historical trajectory. Some recent approaches have utilized predicted (or even actual) future temporal information to enhance location prediction, effectively capturing time-specific user behaviors. For example, some methods explicitly incorporate future timestamps as query inputs to model time-dependent mobility patterns (Feng et al. 2024; Deng et al. 2025b). Alternatively, other techniques (Yang, Liu, and Zhao 2022; Song, Ren, and Lu 2025) jointly optimize timestamp prediction and location forecasting through integrated temporal modeling to improve predictive accuracy. However, there still lacks systematic studies on the key role of future spatiotemporal contexts in location prediction, which, as we show in this paper, are highly informative for the future locations. For example, knowing how much time and distance a user will travel could serve as a strong indicator for forecasting the user’s next location.
@@
Against this background, we propose a novel prediction paradigm by first learning to predict the future spatiotemporal contexts, which are then fed as additional inputs to a base location prediction model to boost the location prediction performance. This paradigm has the following two-fold advantages. On one hand, human mobility patterns exhibit inherent heterogeneity. For instance, entertainment activities and daily routines (e.g., home-workplace commuting) demonstrate distinct spatiotemporal patterns: the former often exhibits more variable spatiotemporal distributions, while the latter tends to be less variable. By modeling future spatiotemporal contexts, we can effectively accommodate such heterogeneity, in particular, benefiting the more variable entertainment activities, where traditional sequence models often fail (as evidenced by our experiments later). On the other hand, the spatiotemporal preferences of users also exhibit different characteristics. For example, two users might have the same lunch schedule, but their travel preferences differ: a convenience-oriented user may take 10 minutes to go to nearby places like an office cafeteria within 500m; in contrast, an adventurous user may be willing to spend half an hour traveling 5 km to go to a fancy restaurant. Subsequently, modeling the future spatiotemporal contexts of users can effectively benefit such a scenario. Our empirical analysis also shows that incorporating the future spatiotemporal contexts can significantly reduce mobility entropy (Zhang, Zhao, and Chen 2022; Sun et al. 2024), thereby showing great potential to improve the predictability of the next location (see Figure 2 in Section Empirical Analysis on Future Spatiotemporal Contexts for details).
@@
  • We evaluate STRelay integrated with four state-of-the-art location prediction base models on four real-world trajectory datasets. Results demonstrate that STRelay consistently improves prediction performance across all cases by 3.19%-11.56%. Additionally, we find that the future spatiotemporal contexts are particularly helpful for entertainment-related locations and also for user groups who prefer traveling longer distances, which is indeed complementary to the base location prediction models that often excel at modeling daily routine patterns.
Related Work
Location prediction is a key problem in human mobility modeling, aiming to predict a user’s future location based on their historical mobility trajectories. Traditional next POI prediction methods focus on learning mobility patterns from user historical trajectories. Early approaches relied on features such as location visit frequency (Noulas et al. 2012) or activity preferences. Markov chain-based models (Rendle, Freudenthaler, and Schmidt-Thieme 2010; Cheng et al. 2013; Feng et al. 2015) capture sequential transitions between locations, while recurrent neural network (RNN)-based models (Liu et al. 2016; Zhao et al. 2019) and attention-based models (Lian et al. 2020) effectively model long-term dependencies in mobility trajectories. The Flashback mechanism (Yang et al. 2020) enhances prediction accuracy by leveraging historical hidden states with similar spatiotemporal contexts, inspiring subsequent works (Cao et al. 2021; Li et al. 2021; Wu et al. 2022; Liu et al. 2022; Rao et al. 2022). Additionally, Graph Neural Networks (GNNs) have been employed to enrich representations by constructing spatiotemporal transition graphs (Rao et al. 2022; Yin et al. 2023), historical distance graphs (Qin et al. 2023), or user preference graphs (Xu et al. 2023). However, these methods typically produce static predictions for all future times based on a given mobility trace, limiting their adaptability to dynamic real-world scenarios. Furthermore, several methods incorporate predicted or even actual future temporal information to guide location predictions, capturing time-specific user behaviors. For example, they utilize future timestamps as query timestamps (Gao, Tang, and Liu 2012; Yang et al. 2019; Li et al. 2021; Feng et al. 2024; Deng et al. 2025b) or a query time interval (Zhao et al. 2019, 2020; Feng et al. 2018) to predict a user’s location after a specified period. Others employ a fixed query time threshold to forecast locations within an upcoming time window (Feng et al. 2015, 2018). An alternative approach integrates timestamp prediction with location forecasting to improve predictive accuracy (Yang, Liu, and Zhao 2022; Sun et al. 2024; Song, Ren, and Lu 2025). However, there still lacks systematic studies on the key role of future spatiotemporal contexts in location prediction, which, as we show in this paper, are highly informative for the future locations.
@@
Empirical Analysis on Future Spatiotemporal Contexts
We conduct an empirical analysis from the perspective of mobility entropy to reveal the importance of the future spatiotemporal context for next location prediction. Specifically, we first focus solely on the trajectory sequence and compute the mobility entropy (Zhang, Zhao, and Chen 2022) for each user as follows:
@@
(a) Tokyo
(b) Singapore
@@
Furthermore, we introduce future temporal context into the mobility entropy through categorizing locations visited by users into their respective future $t$, and then compute the mean mobility entropy after accounting for future temporal context as follows:
@@
The distributions of the mobility entropy on our Tokyo and Singapore datasets (more details below) are shown in Figure 2. We observe that the introduction of temporal context or spatial context can effectively reduce mobility entropy in both datasets. This reduction suggests that involving future spatiotemporal contexts for next location prediction has great potential to reduce the uncertainty of the next location, ultimately resulting in higher predictability.
STRelay
@@
Future Spatiotemporal Context Modeling
@@
Discretization of Future Spatiotemporal Contexts.
@@
where any distances exceeding $N\Delta d$ are capped at $M\Delta t$, and each future spatial distance between consecutive check-ins is assigned to its corresponding bin and represented as an $N$-dimensional one-hot vector $\rho\in\mathbb{R}^{N}$. Afterwards, for all candidate future temporal and spatial intervals, we embed them into a latent space as learnable embeddings $E^{\mathcal{T}}\in\mathbb{R}^{M\times d}$ and $E^{\mathcal{D}}\in\mathbb{R}^{N\times d}$, respectively, where $d$ is the dimension of the embedding space.
@@
Representation of Future Spatiotemporal Contexts.
Human mobility exhibits strong spatiotemporal coupling, where daily routines (e.g., home-work commutes) typically involve both short time intervals and moving distances, while entertainment activities often exhibit longer travel times and distances. To capture such inherent spatiotemporal correlation, we adopt a relaying manner that generates future spatiotemporal context representations by mimicking an individual’s sequential decisions about how much time and how far to move next (Yuan et al. 2022; Deng et al. 2025a). Specifically, we first learn a temporal context representation and subsequently condition on the learnt temporal context representation to output a spatial context representation.
Intuitively, the temporal context for the next check-in depends on the user and the current time. For example, a user at a workplace on a weekday morning may check in at a nearby café within 1 hour, while the same user at a park on a weekend may check in at a distant restaurant after 5 hours. To capture this dependency, when predicting next location $l_{i+1}$, we first embed the user $u$ into the low-dimensional embedding space, denoted as $e^{u}\in\mathbb{R}^{d}$. Next, for the current timestamp $t_{i}$, we convert it to an hour-in-week representation (Deng et al. 2025b) and embed it into a latent space as a learnable embedding $e^{t}_{i}\in\mathbb{R}^{d}$, which can naturally capture temporal regularities between weekdays and weekends. Then, we concatenate them to get the temporal query embedding $e^{\tau}_{i}$ for predicting future temporal context:
@@
Then, we use temporal query embedding to query all of the candidate temporal interval embeddings $E^{\mathcal{T}}$ via an attention mechanism (Vaswani et al. 2017), producing a future temporal context representation $\hat{e}^{\tau}_{i+1}$:
@@
History Encoding and Location Prediction
Base Location Prediction Model for History Encoding. Our STRelay framework is designed to flexibly integrate with different location prediction models under encoder-decoder architectures. While $e^{st}_{i+1}$ captures the future temporal and spatial contexts, to predict the next POI $l_{i+1}$, we also need to encode historical trajectory information. Given a base location prediction model $g$, we input the historical trajectory $X$ to obtain $h_{i}\in\mathbb{R}^{d}$, which encodes the historical sequential spatiotemporal information until $t_{i}$:

### v2

# STRelay: A Universal Spatio-Temporal Relaying Framework for Location Prediction over Human Trajectory Data

@@ Abstract
Next location prediction is a critical task in human mobility modeling, enabling applications like travel planning and urban mobility management. Existing methods mainly rely on historical spatiotemporal trajectory data to train sequence models that directly forecast future locations. However, they often overlook the importance of the future spatiotemporal contexts, which are highly informative for the future locations. For example, knowing how much time and distance a user will travel could serve as a critical clue for predicting the user’s next location. Against this background, we propose STRelay, a universal SpatioTemporal Relaying framework explicitly modeling the future spatiotemporal context given a human trajectory, to boost the performance of different location prediction models. Specifically, STRelay models future spatiotemporal contexts in a relaying manner, which is subsequently integrated with the encoded historical representation from a base location prediction model, enabling multi-task learning by simultaneously predicting the next time interval, next moving distance interval, and finally the next location. We evaluate STRelay integrated with five state-of-the-art location prediction base models on four real-world trajectory datasets. Results demonstrate that STRelay consistently improves prediction performance across all cases by 2.49%-11.30%. Additionally, we find that the future spatiotemporal contexts are particularly helpful for entertainment-related locations and also for user groups who prefer traveling longer distances. The performance gain on such non-daily-routine activities, which often suffer from higher uncertainty, is indeed complementary to the base location prediction models that often excel at modeling regular daily routine patterns.
Index Terms:
Spatiotemporal Data Mining, User mobility, Location prediction
1 Introduction
With the rapid growth of location-based social media, users are increasingly inclined to share their activities on Location-Based Social Networks (LBSNs), generating rich digital footprints for human mobility studies [1]. A key task in human mobility modeling is predicting a user’s next location based on the user’s historical mobility trajectory [2], which is a fundamental building block supporting various location-based applications, such as travel planning [3], taxi ride-sharing scheduling [4], and user behavior patterns analyzing [5].
@@ 1 Introduction
To leverage the spatiotemporal characteristics of trajectory data, various methods have been proposed for next Point of Interest (POI) prediction [^1] Given the context of human trajectories, we do not distinguish the two terms “location” and “POI” throughout this paper.. Traditional next POI prediction methods forecast future locations relying solely on historical spatiotemporal information. For example, sequential transitions are typically captured using Markov chain models [6, 7, 8], while long-term dependencies are modeled through RNNs [9, 10] and attention-based architectures [11, 12, 13, 14]. More recently, some methods using additional context, such as historical spatiotemporal contexts [15] and social networks [16, 17, 18], have been proposed for further performance improvement. However, these methods overlook the time information of the predicted next location, and thus always make the same prediction given a historical trajectory. Some recent approaches have utilized predicted (or even actual) future temporal information to enhance location prediction, effectively capturing time-specific user behaviors. For example, some methods explicitly incorporate future timestamps as query inputs to model time-dependent mobility patterns [19, 20]. Alternatively, other techniques [21, 22] jointly optimize timestamp prediction and location forecasting through integrated temporal modeling to improve predictive accuracy. However, there still lack systematic studies on the key role of future spatiotemporal contexts in location prediction, which, as we show in this paper, are highly informative for the future locations. For example, knowing how much time and distance a user will travel could serve as a strong indicator for forecasting the user’s next location.
@@
Against this background, we propose a novel prediction paradigm by first learning to predict the future spatiotemporal contexts, which are then fed as additional inputs to a base location prediction model to boost the location prediction performance. This paradigm has the following two-fold advantages. On the one hand, human mobility patterns exhibit inherent heterogeneity. For instance, entertainment activities and daily routines (e.g., home-workplace commuting) demonstrate distinct spatiotemporal patterns: the former often exhibits more variable spatiotemporal distributions, while the latter tends to be less variable. By modeling future spatiotemporal contexts, we can effectively accommodate such heterogeneity, in particular, benefiting the more variable entertainment activities, where traditional sequence models often fail (as evidenced by our experiments later). On the other hand, the spatiotemporal preferences of users also exhibit different characteristics. For example, two users might have the same lunch schedule, but their travel preferences differ: a convenience-oriented user may take 10 minutes to go to nearby places like an office cafeteria within 500m; in contrast, an adventurous user may be willing to spend half an hour traveling 5 km to go to a fancy restaurant. Subsequently, modeling the future spatiotemporal contexts of users can effectively benefit such a scenario. Our empirical analysis also shows that incorporating the future spatiotemporal contexts can significantly reduce mobility entropy [23, 24], thereby showing great potential to improve the predictability of the next location (see Figure 2 in Section 3.2 for details).
@@
  • We evaluate STRelay integrated with five state-of-the-art location prediction base models on four real-world trajectory datasets. Results demonstrate that STRelay consistently improves prediction performance across all cases by 2.49%-11.30%. Additionally, we find that the future spatiotemporal contexts are particularly helpful for entertainment-related locations and also for user groups who prefer traveling longer distances, which is indeed complementary to the base location prediction models that often excel at modeling daily routine patterns.
2 Related Work
Predicting an individual’s next location is a fundamental task in human mobility modeling, which aims to predict a user’s future location based on their historical mobility trajectories. Earlier approaches primarily exploit diverse mobility patterns, such as personal activity preferences [25, 26], historical visit frequency [2, 27], social ties [28, 29], and POI or trajectory embeddings learned through graph embedding techniques [30, 31, 32, 33]. Nevertheless, these methods generally fall short in capturing the sequential dynamics of human movement, a factor repeatedly shown to be critical for achieving high-accuracy next-location prediction [9].
To more effectively capture the sequential patterns that traditional methods fail to model, sequence-based approaches shift the focus to individual trajectories, explicitly modeling transition patterns across successive POIs in check-in sequences. Early efforts relied heavily on Markov Chain (MC) models to predict the next POI by estimating location-transition probabilities. For example, FPMC [6] integrates Markov chains with matrix factorization to derive personalized transition matrices; this paradigm was later augmented with spatial constraints [7, 8]. Despite their effectiveness on short-range dependencies, these methods struggle to capture long-term sequential patterns in human mobility trajectories. To address long-range dependencies, Recurrent Neural Networks (RNNs, LSTMs, GRUs) and Transformer architectures [34] have become dominant in next-location prediction. Given that real-world trajectories are typically sparse and irregular, recent models increasingly incorporate spatiotemporal contexts into recurrent units. A widely adopted strategy is to feed temporal and spatial intervals between consecutive check-ins as auxiliary inputs. For instance, Distance2Pre [35] conditions predictions on pairwise geographic distances; STRNN [9] constructs time- and space-specific transition matrices to enrich RNN hidden states; HST-LSTM [36] extends LSTM gates with spatiotemporal distance; STGN [10] capture sequential patterns of users by introducing dedicated spatial and temporal gates into LSTM cells; NeuNext [37] jointly optimizes location-context prediction and next-location prediction; STAN [38] employs spatiotemporal attention to model non-adjacent dependencies; GeoSAN [11] combines hierarchical geographic gridding with self-attention mechanisms, and flashback mechanism [15] considers the two universal human mobility laws, i.e., spatial regularity and temporal periodicity, for location prediction. Specifically, it enhances prediction accuracy by leveraging historical hidden states with similar spatiotemporal contexts, inspiring subsequent works [39, 40, 41, 42, 16].
Beyond recurrent architectures, graph-based methods adopt a global perspective by jointly modeling all users’ trajectories. This strategy captures long-range sequential regularities that are difficult to discover from individual sequences alone and enhances POI representations by incorporating influences from locations outside a user’s current trajectory. For example, STP-UDGAT [43] applies Graph Attention Networks (GAT) on manually constructed spatial-proximity and frequency-transition graphs to capture both local and global POI correlations; SGRec [44] builds dynamic transition graphs by connecting each POI to its successively visited neighbors and employs GAT for representation learning; Graph-Flashback [16] refines POI embeddings via a spatiotemporal knowledge graph; AGRAN [45] jointly optimizes an adaptive graph structure alongside the sequential predictor in an end-to-end manner; SNPM [46] introduces a dynamic neighbor graph and models timestamp-level dependencies to better capture transition patterns; AGCL [47] constructs multiple adaptive graphs (each reflecting a distinct type of POI dependency) and fuses their outputs into a comprehensive multi-faceted POI representation; MobGT [18] simultaneously encodes user mobility preferences and intrinsic location attributes to extract rich spatial-temporal features; GETNext [21] builds global transition graphs from collective mobility patterns and leverages GNNs for POI representation learning; Diff-POI [17] exploits transition and distance graphs to derive spatiotemporal embeddings and introduces a diffusion-based sampling strategy to model spatial preferences; and LoTNext [48] employs a Long-Tailed Graph Adjustment module to mitigate the impact of long-tailed nodes within the User-POI Interaction Graph. MCLP [24] employs Latent Dirichlet Allocation (LDA) to discover latent user preferences and leverages a Transformer architecture to model sequential transition patterns.
More recently, researchers have begun leveraging the powerful reasoning and generation capabilities of Large Language Models (LLMs) for human mobility prediction. For example, LLM-Move [49] and AgentMove [50] reformulate next-location prediction as a text generation task by converting check-in sequences into natural-language prompts and directly querying frozen LLMs such as ChatGPT; LLMob [51] employs an agent-based framework that simulates user motivations and activity patterns to generate synthetic mobility trajectories for different personas; MobilityLLM [52] replaces the conventional Transformer backbone with an LLM and fine-tunes it on multiple mobility analysis tasks; LLM4POI [53] constructs an instruction-tuning dataset from context-enriched mobility records (including POI categories and similar historical sequences) to fine-tune LLMs; GenUP [54] further injects explicit user profiles into system prompts to model individual personality traits; QT-Mob [55] overcomes the limitations of ID-based representations and shallow LLM adaptation, achieving state-of-the-art performance across mobility benchmarks; SILO [56] builds a hybrid semantic space that seamlessly integrates traditional ID embeddings, context-aware semantics, and auxiliary information, thereby enabling joint modeling of sequential patterns and rich contextual knowledge. Note that LLM-based approaches are beyond the scope of this paper, as our focus remains on dedicated neural architectures designed specifically for future spatiotemporal contexts modeling.
Furthermore, several methods incorporate predicted or even actual future temporal information to guide location predictions, capturing time-specific user behaviors. For example, they utilize future timestamps as query timestamps [27, 33, 40, 19, 20] or a query time interval [10, 37, 57] to predict a user’s location after a specified period. Others employ a fixed query time threshold to forecast locations within an upcoming time window [8, 57]. An alternative approach integrates timestamp prediction with location forecasting to improve predictive accuracy [21, 24, 22]. However, there still lack systematic studies on the key role of future spatiotemporal contexts in location prediction, which, as we show in this paper, are highly informative for the future locations.
@@ 3.1 Problem Definition
3.2 Empirical Analysis on Future Spatiotemporal Contexts
We conduct an empirical analysis from the perspective of mobility entropy to reveal the importance of the future spatiotemporal context for next location prediction. Specifically, we first focus solely on the trajectory sequence and compute the mobility entropy [23] for each user as follows:
@@ 3.2 Empirical Analysis on Future Spatiotemporal Contexts
(a) Istanbul
(b) Tokyo
(c) Singapore
(d) Moscow
@@
Furthermore, we introduce future temporal context [24] into the mobility entropy through categorizing locations visited by users into their respective future $t$, and then compute the mean mobility entropy after accounting for future temporal context as follows:
@@
The distributions of the mobility entropy on four datasets are shown in Figure 2. We observe that the introduction of temporal context or spatial context can effectively reduce mobility entropy in both datasets. This reduction suggests that involving future spatiotemporal contexts for next location prediction has great potential to reduce the uncertainty of the next location, ultimately resulting in higher predictability.
4 STRelay
@@
4.1 Future Spatiotemporal Context Modeling
@@
4.1.1 Discretization of Future Spatiotemporal Contexts.
@@
where any distances exceeding $N\Delta d$ are capped at $N\Delta d$, and each future spatial distance between consecutive check-ins is assigned to its corresponding bin and represented as an $N$-dimensional one-hot vector $\rho\in\mathbb{R}^{N}$. Afterwards, for all candidate future temporal and spatial intervals, we embed them into a latent space as learnable embeddings $E^{\mathcal{T}}\in\mathbb{R}^{M\times d}$ and $E^{\mathcal{D}}\in\mathbb{R}^{N\times d}$, respectively, where $d$ is the dimension of the embedding space.
@@
4.1.2 Representation of Future Spatiotemporal Contexts.
Human mobility exhibits strong spatiotemporal coupling, where daily routines (e.g., home-work commutes) typically involve both short time intervals and moving distances, while entertainment activities often exhibit longer travel times and distances. To capture such inherent spatiotemporal correlation, we adopt a relaying manner that generates future spatiotemporal context representations by mimicking an individual’s sequential decisions about how much time and how far to move next [58, 59]. Specifically, we first learn a temporal context representation and subsequently condition on the learnt temporal context representation to output a spatial context representation.
Intuitively, the temporal context for the next check-in depends on the user and the current time. For example, a user at a workplace on a weekday morning may check in at a nearby café within 1 hour, while the same user at a park on a weekend may check in at a distant restaurant after 5 hours. To capture this dependency, when predicting next location $l_{i+1}$, we first embed the user $u$ into the low-dimensional embedding space, denoted as $e^{u}\in\mathbb{R}^{d}$. Next, for the current timestamp $t_{i}$, we convert it to an hour-in-week representation [20] and embed it into a latent space as a learnable embedding $e^{t}_{i}\in\mathbb{R}^{d}$, which can naturally capture temporal regularities between weekdays and weekends. Then, we concatenate them to get the temporal query embedding $e^{\tau}_{i}$ for predicting future temporal context:
@@
Then, we use temporal query embedding to query all of the candidate temporal interval embeddings $E^{\mathcal{T}}$ via an attention mechanism [34], producing a future temporal context representation $\hat{e}^{\tau}_{i+1}$:
@@
4.2 History Encoding and Location Prediction
4.2.1 Base Location Prediction Model for History Encoding
Our STRelay framework is designed to flexibly integrate with different location prediction models under encoder-decoder architectures. While $e^{st}_{i+1}$ captures the future temporal and spatial contexts, to predict the next POI $l_{i+1}$, we also need to encode historical trajectory information. Given a base location prediction model $g$, we input the historical trajectory $X$ to obtain $h_{i}\in\mathbb{R}^{d}$, which encodes the historical sequential spatiotemporal information until $t_{i}$: